# Global Buckley–Leverett for Multicomponent Flow in Fractured Media: Isothermal Equation-of-State Coupling and Dynamic Capillarity

### v1

@@ Abstract
We present an isothermal Global Buckley–Leverett framework for multicomponent, multiphase flow in porous and fractured media that retains the interpretability of classical Buckley–Leverett while incorporating essential physics: equation of state-based phase behavior, multicomponent Maxwell–Stefan diffusion, dynamic capillarity, stress-sensitive permeability, and non-Darcy fracture flow. The formulation yields a single global-pressure equation driving the total Darcy flux and an exact fractional-flow decomposition of phase velocities with buoyancy and capillary drifts; inertial effects enter as per-phase damping that renormalizes mobilities. Crucially, the combination of Maxwell–Stefan diffusion and dynamic capillarity renders transport pseudo-parabolic, resolving the loss of strict hyperbolicity that plagues three-phase Buckley–Leverett and ensuring a well-posed initial-value problem. In practice, each time step solves the scalar global-pressure equation, reconstructs phase fluxes via the split, and advances strictly conservative component balances; axisymmetric (cylindrical) forms for radial injection with vertical buoyancy are provided. The model reduces exactly to classical Buckley–Leverett when added physics are disabled, making it a practical backbone for carbon storage and contaminant transport in fractured, compositionally complex reservoirs.
@@ I Introduction
The Buckley–Leverett (BL) picture remains the simplest and most transparent way to reason about displacement in porous media. In its classical setting—two immiscible, incompressible phases obeying Darcy’s law with fixed PVT and negligible capillarity—the governing transport reduces to a scalar conservation law whose wave structure follows directly from the fractional-flow curve, i.e., the mobility-weighted fraction of the total flux carried by each phase. That blend of physical clarity and analytic control is why the original waterflood analysis by Buckley and Leverett, together with Leverett’s capillary scaling, still anchors pedagogy, verification, and quick-look analysis [Buckley1942, Leverett1941].
@@
Difficulties arise as soon as we leave the two-phase world. For three immiscible phases, the transport system becomes two coupled conservation laws on the saturation simplex, and it is now well established that the model can lose strict hyperbolicity. Eigenvalues coalesce along curves, umbilic points emerge, and parts of the state space effectively become elliptic; in those regions the textbook BL construction no longer selects a unique sequence of shocks and rarefactions unless additional physics is introduced to regularize the equations [ShearerTrangenstein1989, Holden1990, Azevedo2010, Azevedo2014]. An elegant response is the global-pressure reformulation, which aims to decouple a single pressure equation from transport. For two phases the decoupling is fully equivalent; for three phases it is equivalent only when the data satisfy a stringent Total Differential (TD) compatibility among relative permeabilities and capillary pressures across the ternary diagram [Chavent2009, DiChiara2010JCP]. These facts already hint that three-phase BL, taken literally, lacks stabilizing physics.
@@
In unconventional (shale and ultra-tight) settings the missing mechanisms become decisive. The flow is often compositional and strongly compressible; phase changes; and components exchange across phases and with the solid via adsorption/desorption. Those effects lie outside the fixed-property, immiscible BL assumptions and instead call for an equation-of-state (EOS) framework that carries phase split, densities, and viscosities consistently [Coats1980]. When composition varies, diffusion is inherently multicomponent and cross-coupled: Maxwell–Stefan theory provides the thermodynamically consistent description and reduces to scalar Fickian diffusion only in special limits [Krishna2014]. In nanoporous matrices, no-slip assumptions fail: gas slippage and molecule–wall collisions introduce Knudsen contributions, while adsorbed layers support surface diffusion along the solid [Javadpour2009]. In fractures, connectivity and high rates make inertial (non-Darcy/Forchheimer) losses measurable, and both matrix permeability and fracture transmissivity evolve with effective stress [Chen2015Forchheimer, Meng2020]. Finally, capillarity is rate dependent: there is strong theoretical and experimental evidence for dynamic capillary pressure and the role of interfacial area, and—crucially for three-phase transport—those terms supply exactly the regularization that restores uniqueness where BL alone admits non-unique constructions [HassanizadehGray1993, Bottero2011].
@@
This paper assembles a Global Buckley–Leverett for $N$ components and $N_{p}$ phases (GBL-$N$) that: (i) keeps the BL interpretation visible through a phase-flux split; (ii) embeds EOS-consistent phase behavior, multicomponent diffusion, dynamic capillarity, non-Darcy fracture flow, stress sensitivity, and explicit fracture coupling; (iii) explains mathematically why classical three-phase BL can lose strict hyperbolicity and how Maxwell–Stefan diffusion and dynamic capillarity restore well-posedness; (iv) admits an exact global-pressure decoupling under TD (and a natural multi-phase extension, gTD), with a projection-based surrogate otherwise; and (v) reduces cleanly to classical BL when its assumptions hold. The operative system appears in §II.8; the global-pressure structure is detailed in §II.5; the classical limit and well-posedness are summarized in §II.6; and fracture modeling is reviewed in §II.7.
@@ II Global Buckley–Leverett for multicomponent, multiphase flow
We consider isothermal flow (temperature $T$ is fixed). The spatial position is the vector $\mathbf{x}$ (unrelated to the symbols $x_{i\alpha}$, which denote phase mole fractions). The overall (bulk) composition is $z=(z_{1},\ldots,z_{N_{c}})$ with $\sum_{i=1}^{N_{c}}z_{i}=1$. Phases are indexed by $\alpha=1,\dots,N_{p}$ (with only $N_{p}^{\star}\!\leq\!N_{p}$ present at equilibrium). We write mass density $\rho_{\alpha}$ and molar density $c_{\alpha}$ for each phase, linked by
@@ II.1 State, closures, and thermodynamics
We adopt a minimal set of primitive variables in each continuum $\kappa\in\{\mathrm{m},\mathrm{f}\}$: pressure $p$, (fixed) temperature $T$, phase saturations $S_{\alpha}^{\kappa}$ (with $\sum_{\alpha}S_{\alpha}^{\kappa}=1$), and a single overall composition $z=(z_{1},\dots,z_{N_{c}})$ (with $\sum_{i}z_{i}=1$). Using overall composition rather than per-phase compositions is the classical “overall–composition” strategy: $z$ is conserved across control volumes, evolves smoothly through phase transitions (appearance/disappearance), and keeps transport strictly conservative [Coats1980].
@@
Here $\nu_{\alpha}\in[0,1]$ is the phase molar fraction (so $N_{p}^{\star}$ is the number of phases with $\nu_{\alpha}>0$), $x_{i\alpha}$ is the mole fraction of component $i$ in phase $\alpha$, and $\rho_{\alpha},\mu_{\alpha}$ are phase density and viscosity.
@@
This is equivalent to $f_{i}^{\alpha}=f_{i}^{\beta}$” for all present $\alpha,\beta$, but avoids redundant $O(N_{p}^{2})$ pairwise constraints. Together with (3), it defines the flash. For cubic EOS (SRK/PR),
@@
where $\varphi_{i\alpha}$ is the fugacity coefficient (we use $\varphi$ to avoid confusion with porosity $\phi$). For electrolyte/brine phases, an activity–coefficient model may replace $\varphi_{i\alpha}$. The relation $f(p,T,\rho)=\text{const}$ applies to pure fluids; for mixtures the equilibrium condition is (4)—equality of each component’s fugacity across phases [Michelsen1982, MichelsenMollerup2007]. We do not commit to a specific EOS; we only require that the map (2) be smooth and thermodynamically consistent [Coats1980].
@@
$\mathbf{J}_{i\alpha}$ $=-\,c_{\alpha}\sum_{j=1}^{N_{c}}\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}\,\nabla x_{j\alpha}\;-\;\mathbf{J}_{i\alpha}^{\mathrm{Kn/surf}},$ (6)
with $c_{\alpha}$ the phase molar density, $\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}$ a (possibly anisotropic) porous-media effective diffusivity, and $\mathbf{J}_{i\alpha}^{\mathrm{Kn/surf}}$ a calibrated correction for slip/Knudsen/surface diffusion [Javadpour2009]. Mass conservation in each phase imposes
@@
Equivalently, in chemical–potential variables,
$\mathbf{J}_{\alpha}$ $=-\,c_{\alpha}\,\mathbf{D}^{\mathrm{MS}}_{\alpha}\,\boldsymbol{\Gamma}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha},$ (8)
$[\boldsymbol{\Gamma}_{\alpha}]_{ij}$ $=\frac{\partial\mu_{i\alpha}}{\partial\ln x_{j\alpha}},$ (9)
where $\boldsymbol{\Gamma}_{\alpha}$ (thermodynamic-factor matrix) and $\mathbf{D}^{\mathrm{MS}}_{\alpha}$ are symmetric positive definite in stable phases [Krishna2014].
Intrinsic permeability $k^{\kappa}$ and porosity $\phi^{\kappa}$ may evolve with effective stress $\sigma^{\prime}$. We adopt compact log-linear laws
@@
with reference values $k_{0}^{\kappa},\phi_{0}^{\kappa}$ at $\sigma_{0}^{\prime}$ and coefficients $\gamma_{k}^{\kappa},\gamma_{\phi}^{\kappa}\geq 0$ fitted to lab data. The effective stress follows Terzaghi/Biot, e.g. $\sigma^{\prime}=\sigma-b\,p$ with Biot coefficient $b\in[0,1]$. More elaborate poroelastic couplings can replace (11) without affecting the conservative transport structure [Meng2020].
@@
which later provides pseudo-parabolic regularization of transport.
@@
This means: if a phase is present ($\nu_{\alpha}>0$), it is exactly at stability limit ($\psi_{\alpha}=0$); if it is absent ($\nu_{\alpha}=0$) the stability test may be strictly positive ($\psi_{\alpha}>0$). In practice we simply run the stability test and phase split (the “flash”) at each $(p,T,z)$; present phases satisfy $\psi_{\alpha}=0$. We then transport $z$ and recover $\{\nu_{\alpha},x_{i\alpha},\rho_{\alpha},\mu_{\alpha}\}$ by flashing, which preserves conservation and smoothness across phase transitions [Coats1980].
@@ II.2 Component conservation with adsorption and matrix–fracture exchange
$\partial_{t}\Gamma_{i}^{\kappa}=\frac{\Gamma_{i,\mathrm{eq}}^{\kappa}(p,T,z)-\Gamma_{i}^{\kappa}}{\tau_{i}^{\kappa}},$ (15)
with relaxation time $\tau_{i}^{\kappa}\geq 0$. Since $\Gamma_{i}^{\kappa}$ appears only in storage, it alters wave speeds without changing the conservative flux structure.
Intercontinuum transfer is phase specific and must include capillarity and gravity consistently. Define the phase potential
@@
and write the dual–continuum transfer (Warren–Root lineage) as
$\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ $=\omega_{\alpha}\,\frac{k_{\mathrm{int}}}{\bar{\mu}_{\alpha}}\,k_{r\alpha}^{\mathrm{eff}}\;\big(\Phi_{\alpha}^{\mathrm{f}}-\Phi_{\alpha}^{\mathrm{m}}\big),$ (17)
$T_{i}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ $=\sum_{\alpha=1}^{N_{p}}\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}\;\bar{c}_{\alpha}\,\bar{x}_{i\alpha},$ (18)
$(\bar{c}_{\alpha},\bar{x}_{i\alpha})$ $=\text{upwind}\big(\mathrm{sign}\,\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}\big).$ (19)
Using $\Phi_{\alpha}$ guarantees the correct sign when capillarity is strong or gravity opposes the imposed pressure gradient and reduces to a pressure-difference form when $p_{c,\alpha}^{\kappa,\mathrm{dyn}}$ and density variations are negligible. We do not use the global pressure $p_{g}^{\kappa}$ here: $p_{g}^{\kappa}$ is a Darcy-level, mobility-weighted scalar for total flux inside a continuum and is not phase specific.
@@
with $T_{mf}$ the geometric transmissibility and “up” chosen by the sign of $F_{\alpha}^{m\to f}$. Expanding $\Phi_{\alpha}$ recovers the standard pressure, capillary, and gravity jumps used in pEDFM, while strict conservation is preserved by equal-and-opposite interface fluxes.

### v2

@@ Abstract
We present an isothermal Global Buckley–Leverett framework for multicomponent, multiphase flow in porous and fractured media that retains the interpretability of classical Buckley–Leverett while incorporating essential physics: equation of state-based phase behavior, multicomponent Maxwell–Stefan diffusion, dynamic capillarity, stress-sensitive permeability, and non-Darcy fracture flow. The formulation yields a single global-pressure equation driving the total Darcy flux and an exact fractional-flow decomposition of phase velocities with buoyancy and capillary drifts; inertial effects enter as per-phase damping that renormalizes mobilities. Crucially, the combination of Maxwell–Stefan diffusion and dynamic capillarity renders transport pseudo-parabolic, resolving the loss of strict hyperbolicity that plagues three-phase Buckley–Leverett and ensuring a well-posed initial-value problem. In practice, each time step solves the scalar global-pressure equation, reconstructs phase fluxes via the split, and advances strictly conservative component balances; axisymmetric (cylindrical) forms for radial injection with vertical buoyancy are provided. The model reduces exactly to classical Buckley–Leverett when added physics are disabled, making it a practical backbone for carbon storage, geothermal exchange, and contaminant transport in fractured, compositionally complex reservoirs.
@@ I Introduction
The Buckley–Leverett (BL) picture remains the simplest and most transparent way to reason about displacement in porous media. In its classical setting—two immiscible, incompressible phases obeying Darcy’s law with fixed pressure-volume-temperature (PVT) and negligible capillarity—the governing transport reduces to a scalar conservation law whose wave structure follows directly from the fractional-flow curve, i.e., the mobility-weighted fraction of the total flux carried by each phase. That blend of physical clarity and analytic control is why the original waterflood analysis by Buckley and Leverett, together with Leverett’s capillary scaling, still anchors pedagogy, verification, and quick-look analysis [8, 22].
@@
Difficulties arise as soon as we leave the two-phase world. For three immiscible phases, the transport system becomes two coupled conservation laws on the saturation simplex, and it is now well established that the model can lose strict hyperbolicity. Eigenvalues coalesce along curves, umbilic points emerge, and parts of the state space effectively become elliptic; in those regions the textbook BL construction no longer selects a unique sequence of shocks and rarefactions unless additional physics is introduced to regularize the equations [32, 18, 3, 4]. An elegant response is the global-pressure reformulation, which aims to decouple a single pressure equation from transport. For two phases the decoupling is fully equivalent; for three phases it is equivalent only when the data satisfy a stringent Total Differential (TD) compatibility among relative permeabilities and capillary pressures across the ternary diagram [9, 13]. Throughout, ‘phases’ refers to fluid phases; the solid matrix is treated as an immobile skeleton that may store adsorbed mass. These facts already hint that three-phase BL, taken literally, lacks stabilizing physics.
@@
In unconventional (shale and ultra-tight) settings the missing mechanisms become decisive. The flow is often compositional and strongly compressible; phase changes; and components exchange across phases and with the solid via adsorption/desorption. Those effects lie outside the fixed-property, immiscible BL assumptions and instead call for an equation-of-state (EOS) framework that carries phase split, densities, and viscosities consistently [12]. When composition varies, diffusion is inherently multicomponent and cross-coupled: Maxwell–Stefan theory provides the thermodynamically consistent description and reduces to scalar Fickian diffusion only in special limits [21]. In nanoporous matrices, no-slip assumptions fail: gas slippage and molecule–wall collisions introduce Knudsen contributions, while adsorbed layers support surface diffusion along the solid [19]. In fractures, connectivity and high rates make inertial (non-Darcy/Forchheimer) losses measurable, and both matrix permeability and fracture transmissivity evolve with effective stress [10, 24]. Finally, capillarity is rate dependent: there is strong theoretical and experimental evidence for dynamic capillary pressure and the role of interfacial area, and—crucially for three-phase transport—those terms supply exactly the regularization that restores uniqueness where BL alone admits non-unique constructions [17, 7].
@@
This paper assembles a Global Buckley–Leverett for $N$ components and $N_{p}$ phases (GBL-$N$) that: (i) keeps the BL interpretation visible through a phase-flux split; (ii) embeds EOS-consistent phase behavior, multicomponent diffusion, dynamic capillarity, non-Darcy fracture flow, stress sensitivity, and explicit fracture coupling; (iii) explains mathematically why classical three-phase BL can lose strict hyperbolicity and how Maxwell–Stefan diffusion and dynamic capillarity restore well-posedness; (iv) admits an exact global-pressure decoupling under TD (and a natural multi-phase extension, gTD), with a projection-based surrogate otherwise; and (v) reduces cleanly to classical BL when its assumptions hold. The operative system appears in §II.8; the global-pressure structure is detailed in §II.5; the classical limit and well-posedness are summarized in §II.6; and fracture modeling is reviewed in §II.7. The present formulation is a Darcy-scale model built on phase-based superficial velocities, relative permeabilities, and Darcy-type momentum closures (with a Forchheimer-type inertial correction used primarily in fractures). It is intended for regimes where Darcy-scale averaging remains meaningful (connected flow pathways and representative averaging) and where additional physics (EOS phase behavior, multicomponent diffusion, dynamic capillarity, stress sensitivity) is required. We emphasize that multiphase flow may also exhibit nonlinear effective laws associated with pore-scale interface mobilization and intermittent pathway flow, often reported as a power-law relation between flow rate and pressure gradient; this regime is not captured by a per-phase Darcy plus Forchheimer closure [15, 38]. Accordingly, the present work does not claim to model that regime; rather, it provides a conservative thermodynamically consistent backbone into which alternative nonlinear Darcy-scale closures may be inserted when required (see also §II.3).
@@ II Global Buckley–Leverett for multicomponent, multiphase flow
We consider isothermal flow (temperature $T$ is fixed). The spatial position is the vector $\mathbf{x}$ (unrelated to the symbols $x_{i\alpha}$, which denote phase mole fractions). The overall composition is $z=(z_{1},\ldots,z_{N_{c}})$ with $\sum_{i=1}^{N_{c}}z_{i}=1$. Phases are indexed by $\alpha=1,\dots,N_{p}$ (with only $N_{p}^{\star}\!\leq\!N_{p}$ present at equilibrium). We write mass density $\rho_{\alpha}$ and molar density $c_{\alpha}$ for each phase, linked by
@@ II.1 State, closures, and thermodynamics
We adopt a minimal set of primitive variables in each continuum $\kappa\in\{\mathrm{m},\mathrm{f}\}$: pressure $p$, (fixed) temperature $T$, phase saturations $S_{\alpha}^{\kappa}$ (with $\sum_{\alpha}S_{\alpha}^{\kappa}=1$), and a single overall composition $z=(z_{1},\dots,z_{N_{c}})$ (with $\sum_{i}z_{i}=1$). Using overall composition rather than per-phase compositions is the classical “overall–composition” strategy: $z$ is conserved across control volumes, evolves smoothly through phase transitions (appearance/disappearance), and keeps transport strictly conservative [12].
@@
Here $\nu_{\alpha}\in[0,1]$ is the phase molar fraction (so $N_{p}^{\star}$ is the number of phases with $\nu_{\alpha}>0$), $x_{i\alpha}$ is the mole fraction of component $i$ in phase $\alpha$, and $\rho_{\alpha},\mu_{\alpha}$ are phase density and viscosity. While the EOS provides phase split and densities and enables the computation of chemical potentials (or fugacity coefficients) used for phase equilibrium, phase viscosities are obtained from separate transport-property correlations evaluated at the flashed state (e.g., Lohrenz–Bray–Clark and corresponding-states / EOS-based viscosity models)[23, 30, 14, 26, 31, 27].
@@
This is equivalent to $f_{i}^{\alpha}=f_{i}^{\beta}$” for all present $\alpha,\beta$, but avoids redundant $O(N_{p}^{2})$ pairwise constraints. Together with (3), it defines the flash. For cubic EOS Soave–Redlich–Kwong(SRK) / Peng–Robinson (PR),
@@
where $\varphi_{i\alpha}$ is the fugacity coefficient (we use $\varphi$ to avoid confusion with porosity $\phi$). For electrolyte/brine phases, an activity–coefficient model may replace $\varphi_{i\alpha}$. The relation $f(p,T,\rho)=\text{const}$ applies to pure fluids; for mixtures the equilibrium condition is (4)—equality of each component’s fugacity across phases [25, 26]. We do not commit to a specific EOS; we only require that the map (2) be smooth and thermodynamically consistent [12].
@@
$\mathbf{J}_{i\alpha}$ $=-\,c_{\alpha}\sum_{j=1}^{N_{c}}\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}\cdot\nabla x_{j\alpha}\;-\;\mathbf{J}_{i\alpha}^{\mathrm{Kn/surf}},$ (6)
with $c_{\alpha}$ the phase molar density, $\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}$ a (possibly anisotropic) porous-media effective diffusivity, and $\mathbf{J}_{i\alpha}^{\mathrm{Kn/surf}}$ a calibrated correction for slip/Knudsen/surface diffusion [19]. By definition of the Maxwell–Stefan diffusive fluxes relative to the phase-average velocity, the diffusive molar fluxes satisfy
@@
That is, $\mathbf{J}_{i\alpha}$ redistributes species within phase $\alpha$ but does not carry net molar flux of the phase. Phase appearance/disappearance and adsorption/desorption modify storage and advective transport through the flashed state map $(p,T,z)\mapsto(\nu_{\alpha},x_{i\alpha},c_{\alpha},\rho_{\alpha})$, but do not violate the within-phase Maxwell–Stefan constraint.
Equivalently, collecting $\mathbf{J}_{i\alpha}$ into $\mathbf{J}_{\alpha}=(\mathbf{J}_{1\alpha},\ldots,\mathbf{J}_{N_{c}\alpha})^{\top}$, one may write the Maxwell–Stefan closure in compact chemical–potential form as
$\mathbf{J}_{\alpha}$ $=-\,c_{\alpha}\,\mathbf{D}^{\mathrm{MS}}_{\alpha}\,\boldsymbol{\mathcal{G}}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha},$ (8)
$[\boldsymbol{\mathcal{G}}_{\alpha}]_{ij}$ $=\frac{\partial\mu_{i\alpha}}{\partial\ln x_{j\alpha}},$ (9)
where $\boldsymbol{\mathcal{G}}_{\alpha}$ is the thermodynamic-factor matrix and $\mathbf{D}^{\mathrm{MS}}_{\alpha}$ is the (effective) Maxwell–Stefan diffusivity operator; both are symmetric positive definite in stable phases [21]. Here $\nabla\boldsymbol{x}_{\alpha}=(\nabla x_{1\alpha},\ldots,\nabla x_{N_{c}\alpha})^{\top}$, and the product $\mathbf{D}^{\mathrm{MS}}_{\alpha}\,\boldsymbol{\mathcal{G}}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}$ is understood as the standard contraction yielding the vector of spatial molar fluxes. In anisotropic media, $\mathbf{D}^{\mathrm{MS}}_{\alpha}$ denotes a block operator acting on $\nabla\boldsymbol{x}_{\alpha}$; in isotropic form $\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}=D^{\mathrm{MS}}_{ij,\alpha}\mathbf{I}$.
Intrinsic permeability $k^{\kappa}$ and porosity $\phi^{\kappa}$ may evolve with effective stress $\sigma^{\prime}$. Exponential (log-linear) effective-stress relations are widely used as compact empirical fits for permeability (and, when needed, porosity) over practical stress ranges. Such forms are consistent with phenomenological crack-closure models (e.g., Gangi-type) and with laboratory trends in tight and fractured rocks [33, 39, 35, 24]. In this work, we therefore parameterize stress sensitivity via
@@
so that permeability and porosity decrease monotonically with increasing effective stress. Here $(k_{0}^{\kappa},\phi_{0}^{\kappa})$ are reference values at $\sigma_{0}^{\prime}$, and $\gamma_{k}^{\kappa},\gamma_{\phi}^{\kappa}\geq 0$ are calibration coefficients obtained from laboratory measurements. Here $\sigma^{\prime}$ denotes effective stress (e.g., $\sigma^{\prime}=\sigma-bp$ with Biot coefficient $b\in[0,1]$). The effective stress follows Terzaghi/Biot, e.g. $\sigma^{\prime}=\sigma-b\,p$ with Biot coefficient $b\in[0,1]$. More elaborate poroelastic couplings can replace (11) without affecting the conservative transport structure [24].
@@
where $\tau_{\alpha}^{\kappa}(S_{\alpha}^{\kappa})$ ins the dynamic capillarity coefficient (units of $Pa\cdot s$), which later provides pseudo-parabolic regularization of transport.
@@
This means: if a phase is present ($\nu_{\alpha}>0$), it is exactly at stability limit ($\psi_{\alpha}=0$); if it is absent ($\nu_{\alpha}=0$) the stability test may be strictly positive ($\psi_{\alpha}>0$). In practice we simply run the stability test and phase split (the “flash”) at each $(p,T,z)$; present phases satisfy $\psi_{\alpha}=0$. We then transport $z$ and recover $\{\nu_{\alpha},x_{i\alpha},\rho_{\alpha},\mu_{\alpha}\}$ by flashing, which preserves conservation and smoothness across phase transitions [12].
@@ II.2 Component conservation with adsorption and matrix–fracture exchange
$\partial_{t}\Gamma_{i}^{\kappa}=\frac{\Gamma_{i,\mathrm{eq}}^{\kappa}(p,T,z)-\Gamma_{i}^{\kappa}}{t_{i}^{\kappa}},$ (15)
with relaxation time $t_{i}^{\kappa}\geq 0$. Since $\Gamma_{i}^{\kappa}$ appears only in storage, it alters wave speeds without changing the conservative flux structure.
Intercontinuum transfer is phase specific and must include capillarity and gravity consistently. Define the phase potential (phase hydraulic potential)
@@
and model matrix–fracture exchange in a dual-porosity (Warren–Root/Kazemi-type) form as a Darcy-like flux driven by the phase-potential difference [36, 20, 16]:
$\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ $=\omega\,k_{\mathrm{int}}\left(\frac{k_{r\alpha}}{\mu_{\alpha}}\right)^{\mathrm{up}}\Big(\Phi_{\alpha}^{\mathrm{f}}-\Phi_{\alpha}^{\mathrm{m}}\Big),$ (17)
$T_{i}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ $=\sum_{\alpha=1}^{N_{p}}\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}\;c_{\alpha}^{\mathrm{up}}\,x_{i\alpha}^{\mathrm{up}},$ (18)
where $\omega$ is the (geometry-dependent) shape factor (units $1/\mathrm{m}^{2}$) and $k_{\mathrm{int}}$ is an interporosity (matrix-to-fracture) permeability scale. The mobility $\left(k_{r\alpha}/\mu_{\alpha}\right)^{\mathrm{up}}$ and the compositional factors $(c_{\alpha}x_{i\alpha})^{\mathrm{up}}$ are upwinded with respect to the sign of $\mathcal{T}_{\alpha}^{\mathrm{m}\leftrightarrow\mathrm{f}}$. Using $\Phi_{\alpha}$ ensures the correct transfer direction when capillarity is strong and/or gravity opposes the imposed pressure gradient; in the limit of negligible capillarity and density contrasts, (17) reduces to a pressure-difference transfer law. We emphasize that we do not use the global pressure $p_{g}^{\kappa}$ in the transfer term because $p_{g}^{\kappa}$ is a mobility-weighted scalar that drives total flux within each continuum rather than a phase-specific potential.
@@
with $T_{mf}$ the geometric transmissibility and “up” chosen by the sign of $F_{\alpha}^{m\to f}$. Expanding $\Phi_{\alpha}$ recovers the standard pressure, capillary, and gravity jumps used in projection-based embedded discrete fracture model (pEDFM), while strict conservation is preserved by equal-and-opposite interface fluxes.